# Approximating Sparse Matrices and their Functions using Matrix-vector products

### v2

@@ 4.3 Sparse dataset examples
We now illustrate BaMRAM and SpaMRAM using the SuiteSparse Matrix Collection [21], using the four matrices below. The first three matrices are sparse and the last matrix is banded.
@@
For the sparse matrices $\mathtt{west1505}$, $\mathtt{CSphd}$ and $\mathtt{tols2000}$, we use SpaMRAM to recover their matrix exponential. For the banded matrix $\mathtt{mhd3200b}$, we use BaMRAM to recover its matrix exponential, matrix square root and matrix logarithm. The results are illustrated in Figure 4.
@@
Figure 4: The accuracy of BaMRAM (left) and SpaMRAM (right) using sparse matrices from [21].
@@
In Figure 4, we observe that our algorithms are able to approximate $f(\bm{A})$ with good accuracy. In Figure 4a, we notice that the error decreases exponentially until the error stagnates at some point. The stagnation is due to the error incurred in the computation of matrix-vector products $f(\bm{A})\bm{b}$, and can be improved by taking more than 20 Krylov steps. For banded matrices, we consistently obtain exponential convergence rate whose convergence rate is dictated by the smoothness of $f$. In Figure 4b, we note that the error decreases at a rapid speed until it stagnates due to machine precision for $\mathtt{CSphd}$ or slows down due to the sparsity pattern in the case of $\mathtt{west1505}$ and $\mathtt{tols2000}$. As observed in prior experiments, the behaviour is more irregular when the matrix has a more general sparsity pattern.
@@
The approximation to $f(\bm{A})$ in the above experiments can be highly accurate when the bandwidth of the matrix is small or the matrix has a sparsity pattern that makes $f(\bm{A})$ approximately sparse, in the sense that only few entries in each row/column are large in magnitude. Our algorithms are particularly useful when $\bm{A}$ can only be accessed through matrix-vector products and we require an approximation to the entire matrix function. An example is in network analysis where we are given an adjacency matrix $\bm{A}$ and the off-diagonal entries of $e^{\bm{A}}$ describe the communicability between two vertices and the diagonal entries of $e^{\bm{A}}$ describe the subgraph centrality of a vertex [6, 25, 26]. The sparse matrix $\mathtt{CSphd}$ in the above experiment is an adjacency matrix for some underlying graph, and with about $180$ matrix-vector products we can approximate its matrix exponential to $10^{-15}$ accuracy. This means that our algorithm can approximate the communicability between any two vertices and the subgraph centrality of any vertex of $\mathtt{CSphd}$ with $10^{-15}$ accuracy.
@@ 5 Discussion and extensions
In this paper, we presented two algorithms for approximating functions of sparse matrices and matrices with similar structures using matrix-vector products only. We exploited the decay bound from [8] to recover the entries of $f(\bm{A})$ that are large in magnitude. The task of approximating functions of matrices is not the only application for SpaMRAM and BaMRAM. These two algorithms can be used more generally in applications where we are able to perform matrix-vector products with the matrix in question and the matrix we want to approximate satisfies a decay bound similar to (5) or (12). An example is the decay bound for spectral projectors [5], computed, for example by a discretized contour integral. Therefore the algorithms can be extended naturally to other similar applications.
In many applications, we also want to approximate quantities related to $f(\bm{A})$ such as the trace of $f(\bm{A})$ and the Schatten $p$-norm of $f(\bm{A})$, which for the trace of functions of sparse symmetric matrices has been studied in [30]. Since SpaMRAM and BaMRAM provide an approximation to $f(\bm{A})$, we can use $\widehat{f(\bm{A})}$ to estimate the quantity of interest. For example, the trace of a matrix function is an important quantity appearing in network analysis through the Estrada index [26] and statistical learning through the log-determinant estimation [17], among others. Therefore the algorithms can also be used to estimate quantities related to matrix functions.
SpaMRAM uses compressed sensing to recover the matrix function row by row. This algorithm is simple, however we may have access to extra information such as the sparsity pattern of the matrix. This information can potentially be used to speed up the algorithm and increase the accuracy of the algorithm. There are techniques in compressed sensing such as variance density sampling [44], which improves signal reconstruction using prior information. The improvement of SpaMRAM and the use of SpaMRAM and BaMRAM in different applications are left for future work.
@@ Acknowledgements
The authors are grateful to Jared Tanner for his helpful discussions and for pointing out useful literature in compressed sensing. TP is supported by the Heilbronn Institute for Mathematical Research. YN is supported by EPSRC grants EP/Y010086/1 and EP/Y030990/1.

### v3

@@ 4.3 Sparse dataset examples
We now illustrate BaMRAM and SpaMRAM using the SuiteSparse Matrix Collection FloridaDataset2011, using the four matrices below. The first three matrices are sparse and the last matrix is banded.
@@
(a) $\mathtt{west1505}$
(b) $\mathtt{tols2000}$
Figure 4: Eigenvalue spectrum of $\mathtt{west1505}$ and $\mathtt{tols2000}$
@@
The eigenvalue spectrum of $\mathtt{west1505}$ and $\mathtt{tols2000}$ are depicted in Figure 4. For the sparse matrices $\mathtt{west1505}$, $\mathtt{CSphd}$ and $\mathtt{tols2000}$, we use SpaMRAM to recover their matrix exponential. For the banded matrix $\mathtt{mhd3200b}$, we use BaMRAM to recover its matrix exponential, matrix square root and matrix logarithm. The results are illustrated in Figure 5.
@@
Figure 5: The accuracy of BaMRAM (left) and SpaMRAM (right) using sparse matrices from FloridaDataset2011.
@@
In Figure 5, we observe that our algorithms are able to approximate $f(\bm{A})$ with good accuracy. In Figure 5a, we notice that the error decreases exponentially until the error stagnates at some point. The stagnation is due to the error incurred in the computation of matrix-vector products $f(\bm{A})\bm{b}$, and can be improved by taking more than 20 Krylov steps. For banded matrices, we consistently obtain exponential convergence rate whose convergence rate is dictated by the smoothness of $f$. In Figure 5b, we note that the error decreases at a rapid speed until it stagnates due to machine precision for $\mathtt{CSphd}$ or slows down due to the sparsity pattern in the case of $\mathtt{west1505}$ and $\mathtt{tols2000}$. As observed in prior experiments, the behavior is more irregular when the matrix has a more general sparsity pattern.
@@
The approximation to $f(\bm{A})$ in the above experiments can be highly accurate when the bandwidth of the matrix is small or the matrix has a sparsity pattern that makes $f(\bm{A})$ approximately sparse, in the sense that only few entries in each row/column are large in magnitude. Our algorithms are particularly useful when $\bm{A}$ can only be accessed through matrix-vector products and we require an approximation to the entire matrix function. An example is in network analysis where we are given an adjacency matrix $\bm{A}$ and the off-diagonal entries of $e^{\bm{A}}$ describe the communicability between two vertices and the diagonal entries of $e^{\bm{A}}$ describe the subgraph centrality of a vertex BenziEstradaKlymko2013; EstradaHatano2008; EstradaHigham2010. The sparse matrix $\mathtt{CSphd}$ in the above experiment is an adjacency matrix for some underlying graph, and with about $180$ matrix-vector products we can approximate its matrix exponential to $10^{-15}$ accuracy. This corresponds to a total of $180\times 19=3420$ matrix-vector products with $\bm{A}$. This means that our algorithm can approximate the communicability between any two vertices and the subgraph centrality of any vertex of $\mathtt{CSphd}$ with $10^{-15}$ accuracy.
4.4 Scalability test
This section investigate the asymptotic scalability of BaMRAM and SpaMRAM on matrices larger than those considered in the previous sections. We consider the following clases of synthetic matrices, similar to those in Section 4.1:
  • Banded case: $n\times n$ symmetric $2$-banded matrix with its nonzero entries drawn i.i.d. from $\mathcal{N}(0,1)$. The matrix is rescaled so that its $2$-norm equals $0.5$.
  • Sparse case: $n\times n$ symmetric sparse matrix created using the MATLAB command sprandsym with density $\frac{1}{2n}$. This matrix is also rescaled to have $2$-norm equal to $0.5$.
We vary the matrix dimension $n$ from $400$ to $25600$ and record the number of matrix-vector products with $\bm{A}$ required for the methods to meet a prescribed tolerance. The dotted line in the figures corresponds to the number of matrix-vector products with $\bm{A}$ equal to $n$, which is the number required to recover a general dense matrix. Throughout, we use the matrix exponential, where the action $\bm{x}\mapsto f(\bm{A})\bm{x}$ is approximated using $10$ Krylov iterations. All other parameters are identical to those used in the previous experiments.
(a) BaMRAM
(b) SpaMRAM
Figure 6: Scalability of BaMRAM (left) and SpaMRAM (right) for matrices with roughly $\mathcal{O}(1)$ nonzeros per row and column.
The results are shown in Figure 6. Both matrix families have $\mathcal{O}(1)$ nonzero entries per row and column, which suggests that the corresponding matrix exponential may contain only a small number of dominant entries.
In the banded case (Figure 6a), we observe that the number of matrix-vector products with $\bm{A}$ required by BaMRAM is essentially independent of the matrix dimension for all tolerances considered. This suggests that, for matrices with bounded bandwidth, BaMRAM can effectively recover the dominant entries of $f(\bm{A})$ using a number of matrix-vector products with $\bm{A}$ that does not grow with $n$.
For the sparse case, SpaMRAM exhibits a more irregular behavior as the matrix dimension increases, which can be attributed to the unknown sparsity pattern of the original matrix $\bm{A}$. Nevertheless, the number of matrix-vector products grows slowly with $n$ and remains below the dotted line for large $n$ and moderate tolerances. Although the curves exhibit mild fluctuations, the overall scaling is sublinear for this test matrix, indicating that the SpaMRAM effectively exploits the localized structure of the underlying matrix function. In both cases, tighter tolerances lead to an increased number of matrix-vector products, while still maintaining a favorable complexity compared to recovering a general dense matrix.
@@ 5 Discussion and extensions
In this paper, we presented two algorithms for approximating functions of sparse matrices and matrices with similar structures using matrix-vector products only. We exploited the decay bound from Benzi2007 to recover the entries of $f(\bm{A})$ that are large in magnitude. The task of approximating functions of matrices is not the only application for SpaMRAM and BaMRAM. These two algorithms can be used more generally in applications where we are able to perform matrix-vector products with the matrix in question and the matrix we want to approximate satisfies a decay bound similar to (5) or (12). An example is the decay bound for spectral projectors BenziBoitoRazouk2013, computed, for example by a discretized contour integral. Therefore the algorithms can be extended naturally to other similar applications.
In many applications, we also want to approximate quantities related to $f(\bm{A})$ such as the trace of $f(\bm{A})$ and the Schatten $p$-norm of $f(\bm{A})$, which for the trace of functions of sparse symmetric matrices has been studied in FrommerRinelliSchweitzer2023. Since SpaMRAM and BaMRAM provide an approximation to $f(\bm{A})$, we can use $\widehat{f(\bm{A})}$ to estimate the quantity of interest. For example, the trace of a matrix function is an important quantity appearing in network analysis through the Estrada index EstradaHigham2010 and statistical learning through the log-determinant estimation CortinovisKressner2021, among others. Therefore the algorithms can also be used to estimate quantities related to matrix functions.
SpaMRAM uses compressed sensing to recover the matrix function row by row. This algorithm is simple, however we may have access to extra information such as the sparsity pattern of the matrix. This information can potentially be used to speed up the algorithm and increase the accuracy of the algorithm. There are techniques in compressed sensing such as variance density sampling KrahmerWard2014, which improves signal reconstruction using prior information. The improvement of SpaMRAM and the use of SpaMRAM and BaMRAM in different applications are left for future work.
@@ Acknowledgements
The authors are grateful to Jared Tanner for his helpful discussions and for pointing out useful literature in compressed sensing.